# Multi-User Personalisation in Human-Robot Interaction: Resolving Preference Conflicts Using Gradual Argumentation

### v2

@@ Abstract
I Introduction
@@ II Preliminaries
II-A Introduction of Argumentation Frameworks
@@
With the attack relation between arguments, the concept of defence appeared. For example, in Fig. 2a, A attacks B and B attacks C, thus A is defending C since it attacks an argument that attacks C. With those definitions, the semantics for accepting arguments were introduced, being a set of arguments $S\subseteq\mathcal{A}$:
@@
The supporting arguments brought new semantics and definitions of relations, such as direct/indirect attackers or supporters of an argument, the safety sets, which are the ones that do not attack and support the same argument, and new admissible and preferred extensions. In the example from Fig. 2b, the set {A, C} is not safe since the set is attacking and supporting B.
(a) AF with only attacks.
(b) A bipolar AF.
Figure 2: Argumentation Framework Examples. In a), A is attacking B and B is attacking C; therefore, A is defending C. In b), A is attacking B, and C is supporting B.
II-B Quantitative Bipolar Argumentation Frameworks
@@ III Related work
Single-user preferences: Traditional methods for adapting a robot’s behaviour to a user’s preferences or conditions encompass various disciplines within artificial intelligence. Central to this is the concept of user modelling, where the robot builds a representation of the user’s characteristics to inform its actions. This model can be based on a variety of information, including demographic data [Maroto_EAAI23], skill level in a training context [Schadenberg_CSR17], or real-time behavioural cues observed during an interaction [Dell_EAAI22]. Among these methods, Reinforcement Learning (RL) is by far the most used approach in assistive robotics. For instance, Park et al.[Park_aaai19] used RL to maximise children’s learning retention in long-term interactions with a robot teacher, where the reward function is designed to account for both children’s engagement and their learning gains. Similarly, Tsiakas et al.[Tsiakas_MDPI18] proposed a RL framework where the reward function combines explicit feedback, like task performance, with implicit feedback from the user’s observed task engagement. However, designing a reward function that accounts for several heterogeneous parameters can be challenging. Therefore, other studies propose using Inverse Reinforcement Learning (IRL), which infers the underlying reward function from expert demonstrations rather than requiring it to be specified manually. Following this principle, Sugiyama et al.[Sugiyama_interspeech12] learn the appropriate reward function by modelling a user’s pairwise decisions, effectively inferring preferences from their choices between two options. More recently, Andriella et al.[Andriella_UMUAI23] introduced the CARESSER framework, designed to actively learn personalised robot-assistive behaviour for cognitive training therapy. Here, the reward function was learnt by combining a therapist’s expertise and their demonstrations, after which a unique policy was estimated using classical RL approaches. Beyond learning-based methods, classical AI planning has also been employed to achieve robot adaptiveness. This approach typically involves creating a symbolic representation of the user’s needs, which is then integrated directly into the planning domain. A recent work by De Benedictis et al.[Benedictis_UMUAI23] developed a system for personalising cognitive stimulation for older adults. Here, the users’ needs, capabilities, and preferences were directly represented in the planning knowledge base. Canal et al.[Canal_AR18] proposed a symbolic high-level task planning for assistive shoe dressing, in which the user’s preferences were gathered from answering simple questions and integrated into the planning domain to generate appropriate assistance. A significant and common limitation of the aforementioned approaches is their foundation in a single-user paradigm. Whether it is a manually engineered reward function for RL, a set of expert demonstrations for IRL, or a coherent set of preferences for a planner, these methods are all designed to create a unified model tailored to an individual user. This paradigm is fundamentally challenged in multi-user scenarios where conflicting preferences make decisions more difficult, especially because preferences are often subjective for each user and challenging to quantify.
@@
IV Multiple users’ preference selection system
This section introduces the modules (see Fig. 1) of our framework used to select a preference in a scenario with multiple users with possibly competing preferences for a specific action. Firstly, the robot observations and their history generate the task and users’ arguments in the arguments module (see Sec. IV-A). Next, these arguments are fed into the preference selector module formalised as a QBAF to determine the selected preference (see Sec. IV-B). Finally, the selected preference is input to the action sequence executor module to parameterise the action. It is important to note that developing a decision-making system to select the best action in a given state is outside the scope of this work.
@@ IV Multiple users’ preference selection system
IV-A Robot observations and arguments modules
@@
IV-B Preference selector module
This module serves as the core computational engine for resolving conflicts and determining the robot’s course of action. Upon receiving the users’ preferences and arguments, the module initiates a four-stage process. Firstly, it analyses the preferences to detect any conflicts (see Sec. IV-B1). Second, it formalises the conflict using our proposed AF (see Sec. IV-B2). Third, a dedicated algorithm processes this framework to select a winning preference and decide the parameters of the robot’s action (see Sec. IV-B3). Finally, we demonstrate that this decision-making process adheres to several properties that ensure its responsiveness in dynamic contexts (see Sec. IV-B4).
IV-B1 Preferences and conflict types
@@
    $\exists o\in O.\;\exists U^{\prime}\subseteq U.\;|U^{\prime}|\geq 1\land\forall i\in U^{\prime}.\;(o,+)\in\hat{p}(i)$
    $\land\forall o^{\prime}\in O.\;o\neq o^{\prime}\to(o^{\prime},+)\notin\hat{p}(i)$
    $\land\big(\forall j\in U\setminus U^{\prime}.\;(o,0)\in\hat{p}(j)\land(o^{\prime},+)\notin\hat{p}(j)\big)$
@@
    $\forall i\in U.\ \big(\forall o\in O.\ (o,+)\notin P_{i}\land\exists o^{\prime}\in O.\ (o^{\prime},0)\in\hat{p}(i)\big)$
@@
    $\exists i,j\in U.\;\exists o,o^{\prime}\in O.\;o^{\prime}\neq o\land i\neq j$
    $\land(o,+)\in\hat{p}(i)\land(o^{\prime},+)\in\hat{p}(j)$
    $\land(o,+)\notin\hat{p}(j)\land(o^{\prime},+)\notin\hat{p}(i)$
@@
    $\exists i,j\in U.\ \exists\ o,o^{\prime}\in O.\ o^{\prime}\neq o\land i\neq j$
    $\land\ (o,-)\in\hat{p}(i)\ \land\ (o^{\prime},-)\in\hat{p}(j)$
    $\land\ (o,-)\notin\hat{p}(j)\ \land\ (o^{\prime},-)\notin\hat{p}(i)$
@@
IV-B2 The MUP-QBAF framework
The arguments received from the arguments module, and defined in the previous section (see Sec. IV-A), are used to construct a Multi-User Preferences Quantitative Bipolar Argumentation Framework (MUP-QBAF):
@@ IV-B2 The MUP-QBAF framework
IV-B3 Multi-user preference selector algorithm
@@
The first step of the algorithm is to calculate the final strengths of the MUP-QBAF’s arguments (line 1). The next step consists of checking if any of the conditions for conflicting scenarios from Sec. IV-B1 are satisfied (line 2). In case of conflict, the algorithm will select an option $o^{*}\in O^{\prime}$ by running $g$ (line 3), which is a function that selects among the set of options $O^{\prime}\subseteq O$ with the highest strength: $o^{*}=g\{o|\sigma(o)=\max_{o^{\prime}\in O}\sigma(o^{\prime}),\mathcal{M},\nu\}$. Those options with the maximum strength are input into the $g$ function, which selects an option given some criteria. If there is only one option in $g$, the conflict is considered resolved, and the selected option is that one.
@@
IV-B4 Properties of MUP-QBAFs dynamic adaptation
@@
V Using MUP-QBAF in robotic frailty assessment
@@
Firstly, the use case, the role of the robot, and the possible conflicts are explained (see Sec. V-A). Then, an example of a possible AF (see Sec. V-B) is provided. Moreover, some examples related to adaptation to dynamic context, which is crucial in HRI, are presented (see Sec. V-C). Finally, an analysis of the proposed AF in different scenarios, and the sensitivity of the arguments’ base score is shown (see Sec. V-D).
V-A Clinical scenario and motivation
@@
V-B Constructing the Argumentation Framework
@@
V-C Dynamic adaptation examples
A crucial requirement in HRI is their rapid adaptation to dynamic changes in the environment and user preferences. Here, we provide with two examples showing the framework’s adaptation given some changes in the AF.
@@ V-C Dynamic adaptation examples
V-D Framework evaluation and sensitivity analysis
@@
TABLE I: Distribution of preference selections for Repeat (R), No Repeat (N-R) and Tie across different argument configurations for the two conflicting scenarios. The first and fourth rows correspond to all the possible argument combinations. The second and fifth rows are the combinations where the argument T1 is activated. The third and last rows are the combinations when the risk of falling arguments (the upper part of the AF) are removed. All the arguments’ base score is 0.5. R means Repeat, and N-R means No Repeat.
Table I presents the selection percentages for various combinations of activated and deactivated arguments across both scenarios. The first row lists all possible combinations, the second row focuses on instances when the risk of falling argument (T1) is activated, and the third row shows the distributions when the risk of falling arguments are not included. In this last scenario, while the argument CR1 remains in the framework, it does not influence the final strengths because there is no path connecting it to the Option Arguments. From this table, we can observe the impact of the risk of falling argument (T1), which skews the distributions towards “no repeat” thereby making the assessment safer based on actual observations of the robot. This effect is particularly noticeable when comparing the values between the first and second rows for each scenario. In contrast, the rows that exclude the risk of falling arguments show a more balanced distribution between repeating and not repeating. This balance is expected since the Option Arguments have an equal number of attackers and supporters in both scenarios. The differences observed in the last row regarding repetition are attributable to the indirect attackers and supporters of the Option Arguments.
Additionally, it is possible to compute the influence of the relations between arguments on the Option Arguments, based on the approach from [Yin_ijcai24] named Relation Attribution Explanations (RAEs). This analysis shows the influence of each relation on the Option Arguments. Table II shows an example of influences for scenario (2), since it contains more arguments and is more complete, with all the arguments activated and base scores at 0.5.
@@
TABLE II: Scenario 2 (caregiver for repeat, care recipient against repeat). Approximate influence of each relation between arguments on the final Option Arguments when all arguments are active and with a base score of 0.5. The value 0 is for strict zeros, and 0.00 refers to values <5<5 · $10^{-3}$.
@@
(a) Mean and standard deviation (95%) distribution of the difference between the final strengths of repeat and no repeat, ($\sigma(r)-\sigma(\neg r)$), for all the combinations in scenario (1). The top plot corresponds to all combinations (activated and deactivated), the second plot corresponds to combinations where T1 is activated, and the third one to combinations where T1 is deactivated.
(b) Percentages of the decisions, in scenario (1), where the caregiver is against repeat. The top plot shows all combinations. At the bottom, combinations where the risk of falling argument (T1) is active. The blue dashed section refers to percentages where the robot decided to repeat the tests; the yellow dashed section refers to not repeating; the green area refers to ties.
@@
Finally, we conducted a sensitivity analysis to evaluate how the base score of a critical argument influences the final decision. We varied the base score of the “risk of falling” argument (T1), which acts as a parameter to make the robot’s decision-making more or less conservative. Fig. 6a illustrates the results, showing the difference in final strength between the repeat and no repeat options ($\sigma(r)-\sigma(\neg r)$) as the base score of T1 is increased from 0 to 1 (all other arguments were held at a base score of 0.5). The plot shows a clear trend: as the T1 base score increases, the strength difference in favour of repeating diminishes. This directly impacts the robot’s choices, as shown in Fig. 6b. The selection percentage for repeat decreases as the T1 base score grows. This confirms that a higher base score for the “risk of falling” argument makes the robot more cautious, eventually reaching a point where it will strictly select the no repeat option when this risk is present.
VI Limitations and outlook
@@
Regarding the AF design, the following limitations are noted: firstly, this work does not implement Argument Mining, which is the field in CA that generates and relates arguments from user input. The arguments in the AF presented in this work were decided through co-design with healthcare professionals. However, in a complete and fully autonomous real-world deployment, Argument Mining will be necessary for getting the users and task arguments. This will become an important challenge for future work, which can be potentially overcome with the usage of LLM [Gorur_arxiv24]. Secondly, all the base scores in Sec. V are set to 0.5, except in the sensitivity analysis. In some scenarios, it may seem reasonable for the system to assign different base scores to users’ arguments. For instance, users might give arguments that are considered very relevant and important to them [Battaglia_icsum24]. Those arguments could have a greater base score. Another factor considered when setting the base scores can be the expertise of the users in the context, e.g. in assistive scenarios, health-related arguments given by the caregiver may be more grounded than those given by the care recipient. While this can improve decision quality, it also raises a risk of systemic bias, prioritising caregiver preferences at the expense of the care recipient’s autonomy. Thirdly, while this study assumes distinct arguments for each user, scenarios where users share arguments but assign different base scores require further exploration. Methods to aggregate individual AFs and base scores into a general AF should be further explored [Dickie_AAMAS25].
Respecting the system’s interaction limitations, we mention the following: firstly, no explanation is generated regarding the reason for the robot’s selection. Explanations could be based on the influence of each argument or their relationships to the final decision (see Sec. V-D). Similarly, if a user complains or questions the reason behind the robot’s decision, the robot must provide a clear and understandable explanation to facilitate the necessary feedback for adjusting its decision. To achieve this, it is essential to identify the necessary and sufficient changes and explanations required within the framework  [Kampik_IJAR24]. Secondly, this work is presented in a particular assistive use case. Even though it is generalisable for any scenario where multiple users have competing preferences, it should be tested and studied in different use cases and in a real-world scenario. Additionally, this work is based on a single-interaction approach, in which a care recipient interacts with the robot once, only to perform the tests. More investigation on how this system behaves in long-term interactions is needed.

### v3

@@ Abstract
keywords:
Assistive Robots, Multi-User Personalisation, Gradual Argumentation, Conflict Resolution
1 Introduction
@@ 2 Preliminaries
2.1 Introduction of Argumentation Frameworks
@@
With the attack relation between arguments, the concept of defence appeared. For example, in Fig. 2(a), A attacks B and B attacks C, thus A is defending C since it attacks an argument that attacks C. With those definitions, the semantics for accepting arguments were introduced, being a set of arguments $S\subseteq\mathcal{A}$:
@@
The supporting arguments brought new semantics and definitions of relations, such as direct/indirect attackers or supporters of an argument, the safety sets, which are the ones that do not attack and support the same argument, and new admissible and preferred extensions. In the example from Fig. 2(b), the set {A, C} is not safe since the set is attacking and supporting B.
(a) AF with A attacking B and B attacking C.
(b) AF with A attacking B and C supporting B.
Figure 2: Argumentation Framework Examples.
2.2 Quantitative Bipolar Argumentation Frameworks
@@ 3 Related work
Single-user preferences: Traditional methods for adapting a robot’s behaviour to a user’s preferences or conditions encompass various disciplines within artificial intelligence. Central to this is the concept of user modelling, where the robot builds a representation of the user’s characteristics to inform its actions. This model can be based on a variety of information, including demographic data [Maroto_EAAI23], skill level in a training context [Schadenberg_CSR17], or real-time behavioural cues observed during an interaction [Dell_EAAI22]. Among these methods, Reinforcement Learning (RL) is by far the most used approach in assistive robotics. For instance, Park et al.[Park_aaai19] used RL to maximise children’s learning retention in long-term interactions with a robot teacher, where the reward function is designed to account for both children’s engagement and their learning gains. Similarly, Tsiakas et al.[Tsiakas_MDPI18] proposed an RL framework where the reward function combines explicit feedback, like task performance, with implicit feedback from the user’s observed task engagement. However, designing a reward function that accounts for several heterogeneous parameters can be challenging. Therefore, other studies propose using Inverse Reinforcement Learning (IRL), which infers the underlying reward function from expert demonstrations rather than requiring it to be specified manually. Following this principle, Sugiyama et al.[Sugiyama_interspeech12] learn the appropriate reward function by modelling a user’s pairwise decisions, effectively inferring preferences from their choices between two options. More recently, Andriella et al.[Andriella_UMUAI23] introduced the CARESSER framework, designed to actively learn personalised robot-assistive behaviour for cognitive training therapy. Here, the reward function was learnt by combining a therapist’s expertise and their demonstrations, after which a unique policy was estimated using classical RL approaches. Beyond learning-based methods, classical AI planning has also been employed to achieve robot adaptiveness. This approach typically involves creating a symbolic representation of the user’s needs, which is then integrated directly into the planning domain. A recent work by De Benedictis et al.[Benedictis_UMUAI23] developed a system for personalising cognitive stimulation for older adults. Here, the users’ needs, capabilities, and preferences were directly represented in the planning knowledge base. Canal et al.[Canal_AR18] proposed a symbolic high-level task planning for assistive shoe dressing, in which the user’s preferences were gathered from answering simple questions and integrated into the planning domain to generate appropriate assistance. A significant and common limitation of the aforementioned approaches is their foundation in a single-user paradigm. Whether it is a manually engineered reward function for RL, a set of expert demonstrations for IRL, or a coherent set of preferences for a planner, these methods are all designed to create a unified model tailored to an individual user. This paradigm is fundamentally challenged in multi-user scenarios where conflicting preferences make decisions more difficult, especially because preferences are often subjective for each user and challenging to quantify.
@@
4 Multiple users’ preference selection system
This section introduces the modules (see Fig. 1) of our framework used to select a preference in a scenario with multiple users with possibly competing preferences for a specific action. Firstly, the robot observations and their history generate the task and users’ arguments in the arguments module (see Sec. 4.1). Next, these arguments are fed into the preference selector module formalised as a QBAF to determine the selected preference (see Sec. 4.2). Finally, the selected preference is input to the action sequence executor module to parameterise the action. It is important to note that developing a decision-making system to select the best action in a given state is outside the scope of this work.
@@ 4 Multiple users’ preference selection system
4.1 Robot observations and arguments modules
@@
4.2 Preference selector module
This module serves as the core computational engine for resolving conflicts and determining the robot’s course of action. Upon receiving the users’ preferences and arguments, the module initiates a four-stage process. Firstly, it analyses the preferences to detect any conflicts (see Sec. 4.2.1). Second, it formalises the conflict using our proposed AF (see Sec. 4.2.2). Third, a dedicated algorithm processes this framework to select a winning preference and decide the parameters of the robot’s action (see Sec. 4.2.3). Finally, we demonstrate that this decision-making process adheres to several properties that ensure its responsiveness in dynamic contexts (see Sec. 4.2.4).
4.2.1 Preferences and conflict types
@@
    $\exists o\in O.\ \exists U^{\prime}\subseteq U.\ |U^{\prime}|\geq 1\land\forall i\in U^{\prime}.\ (o,+)\in\hat{p}(i)\land\forall o^{\prime}\in O.\ o\neq o^{\prime}\to$
    $(o^{\prime},+)\notin\hat{p}(i)\land\big(\forall j\in U\setminus U^{\prime}.\ (o,0)\in\hat{p}(j)\land(o^{\prime},+)\notin\hat{p}(j)\big)$
@@
    $\forall i\in U.\ \big(\forall o\in O.\ (o,+)\notin P_{i}\land\exists o^{\prime}\in O.\ (o^{\prime},0)\in\hat{p}(i)\big).$
@@
    $\begin{split}\exists\ i,j\in U.\ \exists\ o,o^{\prime}\in O.\ o^{\prime}\neq o\ \land\ i\neq j\ \land\ (o,+)\in\hat{p}(i)\ \land\ (o^{\prime},+)\in\hat{p}(j)\\ \land\ (o,+)\notin\hat{p}(j)\ \land\ (o^{\prime},+)\notin\hat{p}(i)\end{split}$
@@
    $\begin{split}\exists\ i,j\in U.\ \exists\ o,o^{\prime}\in O.\ o^{\prime}\neq o\ \land\ i\neq j\ \land\ (o,-)\in\hat{p}(i)\ \land\ (o^{\prime},-)\in\hat{p}(j)\\ \land\ (o,-)\notin\hat{p}(j)\ \land\ (o^{\prime},-)\notin\hat{p}(i).\end{split}$
@@
4.2.2 The MUP-QBAF framework
The arguments received from the arguments module, and defined in the previous section (see Sec. 4.1), are used to construct a Multi-User Preferences Quantitative Bipolar Argumentation Framework (MUP-QBAF):
@@ 4.2.2 The MUP-QBAF framework
4.2.3 Multi-user preference selector algorithm
@@
The first step of the algorithm is to calculate the final strengths of the MUP-QBAF’s arguments (line 1). The next step consists of checking if any of the conditions for conflicting scenarios from Sec. 4.2.1 are satisfied (line 2). In case of conflict, the algorithm will select an option $o^{*}\in O^{\prime}$ by running $g$ (line 3), which is a function that selects among the set of options $O^{\prime}\subseteq O$ with the highest strength: $o^{*}=g\{o|\sigma(o)=\max_{o^{\prime}\in O}\sigma(o^{\prime}),\mathcal{M},\nu\}$. Those options with the maximum strength are input into the $g$ function, which selects an option given some criteria. If there is only one option in $g$, the conflict is considered resolved, and the selected option is that one.
@@
4.2.4 Properties of MUP-QBAFs dynamic adaptation
@@
5 Using MUP-QBAF in robotic frailty assessment
@@
Firstly, the use case, the role of the robot, and the possible conflicts are explained (see Sec. 5.1). Then, an example of a possible AF (see Sec. 5.2) is provided. Moreover, some examples related to adaptation to dynamic context, which is crucial in HRI, are presented (see Sec. 5.3). Finally, an analysis of the proposed AF in different scenarios, and the sensitivity of the arguments’ base score is shown (see Sec. 5.4).
5.1 Clinical scenario and motivation
@@
5.2 Constructing the Argumentation Framework
@@
5.3 Dynamic adaptation examples
A crucial requirement in HRI is their rapid adaptation to dynamic changes in the environment and user preferences. Here, we provide two examples illustrating the framework’s adaptation in response to certain changes in the AF.
@@ 5.3 Dynamic adaptation examples
5.4 Framework evaluation and sensitivity analysis
@@
Table 1: Distribution of preference selections for Repeat (R), No Repeat (N-R) and Tie across different argument configurations for the two conflicting scenarios. The first and fourth rows correspond to all the possible argument combinations. The second and fifth rows are the combinations where the argument T1 is activated. The third and last rows are the combinations when the risk of falling arguments (the upper part of the AF) are removed. All the arguments’ base score is 0.5. R means Repeat, and N-R means No Repeat.
Table 1 presents the selection percentages for various combinations of activated and deactivated arguments across both scenarios. The first row lists all possible combinations, the second row focuses on instances when the risk of falling argument (T1) is activated, and the third row shows the distributions when the risk of falling arguments are not included. In this last scenario, while the argument CR1 remains in the framework, it does not influence the final strengths because there is no path connecting it to the Option Arguments. From this table, we can observe the impact of the risk of falling argument (T1), which skews the distributions towards “no repeat” thereby making the assessment safer based on actual observations of the robot. This effect is particularly noticeable when comparing the values between the first and second rows for each scenario. In contrast, the rows that exclude the risk of falling arguments show a more balanced distribution between repeating and not repeating. This balance is expected since the Option Arguments have an equal number of attackers and supporters in both scenarios. The differences observed in the last row regarding repetition are attributable to the indirect attackers and supporters of the Option Arguments.
Additionally, it is possible to compute the influence of the relations between arguments on the Option Arguments, based on the approach from [Yin_ijcai24] named Relation Attribution Explanations (RAEs). This analysis shows the influence of each relation on the Option Arguments. Table 2 shows an example of influences for scenario (2), since it contains more arguments and is more complete, with all the arguments activated and base scores at 0.5.
@@
Table 2: Scenario 2 (caregiver for repeat, care recipient against repeat). Approximate influence of each relation between arguments on the final Option Arguments when all arguments are active and with a base score of 0.5. The value 0 is for strict zeros, and 0.00 refers to values <5<5 · $10^{-3}$.
@@
(a) Mean and standard deviation (95%) distribution of the difference between the final strengths of repeat and no repeat, ($\sigma(r)-\sigma(\neg r)$), for all the combinations in scenario (1). The top plot corresponds to the totality of the combinations (activated and deactivated), the second plot corresponds to the combinations where T1 is activated, and the third one corresponds to the combinations where T1 is deactivated.
(b) Percentages of the decisions, in scenario (1), where the caregiver is against repeat. The distribution of all combinations is plotted at the top. At the bot, the combinations where the risk of falling argument is active (T1). The blue dashed section refers to the percentages that the robot decided to repeat the tests. The yellow dashed section refers to not repeating. The green area refers to ties.
@@
Finally, we conducted a sensitivity analysis to evaluate how the base score of a critical argument influences the final decision. We varied the base score of the “risk of falling” argument (T1), which acts as a parameter to make the robot’s decision-making more or less conservative. Fig. 6(a) illustrates the results, showing the difference in final strength between the repeat and no repeat options ($\sigma(r)-\sigma(\neg r)$) as the base score of T1 is increased from 0 to 1 (all other arguments were held at a base score of 0.5). The plot shows a clear trend: as the T1 base score increases, the strength difference in favour of repeating diminishes. This directly impacts the robot’s choices, as shown in Fig. 6(b). The selection percentage for repeat decreases as the T1 base score grows. This confirms that a higher base score for the “risk of falling” argument makes the robot more cautious, eventually reaching a point where it will strictly select the no repeat option when this risk is present.
6 Limitations and outlook
@@
Regarding the AF design, the following limitations are noted: firstly, this work does not implement Argument Mining, which is the field in CA that generates and relates arguments from user input. The arguments in the AF presented in this work were decided through co-design with healthcare professionals. However, in a complete and fully autonomous real-world deployment, Argument Mining will be necessary for getting the users and task arguments. This will become an important challenge for future work, which can be potentially overcome with the usage of LLM [Gorur_arxiv24]. Secondly, all the base scores in Sec. 5 are set to 0.5, except in the sensitivity analysis. In some scenarios, it may seem reasonable for the system to assign different base scores to users’ arguments. For instance, users might give arguments that are considered very relevant and important to them [Battaglia_icsum24, Civit_aamas26]. Those arguments could have a greater base score. Another factor considered when setting the base scores can be the expertise of the users in the context, e.g. in assistive scenarios, health-related arguments given by the caregiver may be more grounded than those given by the care recipient. While this can improve decision quality, it also raises a risk of systemic bias, prioritising caregiver preferences at the expense of the care recipient’s autonomy. Thirdly, while this study assumes distinct arguments for each user, scenarios where users share arguments but assign different base scores require further exploration. Methods to aggregate individual AFs and base scores into a general AF should be further explored [Dickie_AAMAS25].
Respecting the system’s interaction limitations, we mention the following: firstly, no explanation is generated regarding the reason for the robot’s selection. Explanations could be based on the influence of each argument or their relationships to the final decision (see Sec. 5.4). Similarly, if a user complains or questions the reason behind the robot’s decision, the robot must provide a clear and understandable explanation to facilitate the necessary feedback for adjusting its decision. To achieve this, it is essential to identify the necessary and sufficient changes and explanations required within the framework  [Kampik_IJAR24]. Secondly, this work is presented in a particular assistive use case. Even though it is generalisable for any scenario where multiple users have competing preferences, it should be tested and studied in different use cases and in a real-world scenario. Additionally, this work is based on a single-interaction approach, in which a care recipient interacts with the robot once, only to perform the tests. More investigation on how this system behaves in long-term interactions is needed.